# NashOpt - A Python Library for Computing Generalized Nash Equilibria

### v1

@@ 5.1 Game-theoretic linear quadratic regulation
We consider a non-cooperative multi-agent linear-quadratic regulator (LQR) problem where $N=10$ agents control a shared unstable discrete-time linear dynamical system
@@
where $K_{i}\in\mathbb{R}^{n_{i}\times n_{x}}$ is the LQR gain corresponding to weight matrices $Q_{i}\in\mathbb{R}^{n_{x}\times n_{x}}\succeq 0$, such as $Q_{i}=C^{\top}Q_{yi}C$ with $Q_{yi}\succeq 0$, and $R_{i}\in\mathbb{R}^{n_{i}}$, $R_{i}\succ 0$. We approximate $K_{i}$ as the solution of the finite-horizon problem
$K_{i}^{\textrm{L}QR}(K_{-i})\in\arg\min$ $\sum_{k=0}^{N_{\textrm{L}QR}}\left(x_{k}^{\top}Q_{i}x_{k}+u_{i,k}^{\top}R_{i}u_{i,k}\right)$ (17)
$\mathop{\rm s.t.}\nolimits$ $x_{k+1}=(A-B_{-i}K_{-i})x_{k}+B_{i}u_{i,k}.$
We define each agent’s cost function as the squared Frobenius norm of the deviation from its optimal gain $K_{i}^{\textrm{L}QR}(K_{-i})$ in (17):
$f_{i}(K_{i},K_{-i})=\|K_{i}^{\textrm{L}QR}(K_{-i})-K_{i}\|_{F}^{2}.$
@@
The equilibrium is computed as the solution of a Nash equilibrium problem, using the centralized LQR solution $K_{\rm LQR}$ associated with $(A,B,\sum_{i}Q_{i},\mathrm{blkdiag}(R_{1},\dots,R_{N}))$ as initial guess. The problem is solved using the nonlinear least-squares approach (6) (without parameter $p_{0}$).
@@ 5.2 Game-theoretic model predictive control
In (18), $Q_{i}\succeq 0$, $Q_{\Delta u,i}\succeq 0$, and $\epsilon_{i}\geq 0$ is a slack variable used to soften shared output constraints (with linear penalty $q_{\epsilon,i}\geq 0$). For example, if an agent is interested in controlling only a subset $I_{i}$ of the components of the output vector with unit weights, we have $Q_{i}=\sum_{j\in I_{i}}e_{j}e_{j}^{\top}$.
The predicted trajectories $\{x_{k},y_{k},u_{k}\}$ in (18) satisfy (16) over the prediction horizon $k=0,\ldots,T$, with initial condition $x_{0}=x(t)$ and $u_{-1}=u(t-1)$. Note that each agent optimizes the sequence of input increments $\Delta u_{i}$ and the slack variable $\epsilon_{i}$, so that constraints on inputs, input increments, and slacks are local to each agent, while output constraints are shared among all agents, each one softening them with possibly different slacks $\epsilon_{i}$. The number of constraints in each agent’s problem (18) can be reduced by imposing the constraints only on a shorter constraint horizon of $T_{c}<T$ steps. This can be especially useful when $T$ is large but we want to limit the number of Lagrange multipliers and binary variables involved in the MILP reformulation (14) of the GNEP, in which we remove $p$, $\sigma$, and the objective function (12a), as we are only interested in computing the GNE of the noncooperative MPC problem.
The resulting game-theoretic model predictive control (MPC) problem is the GNE problem of finding sequences $\{\Delta u_{i}^{\star}\}_{i=1}^{N}$ (and slacks $\{\epsilon_{i}^{\star}\}_{i=1}^{N}$) optimizing (18) for all agents $i=1,\ldots,N$ simultaneously. In a receding-horizon fashion, only the first inputs $u_{i}(t)=u_{i}(t-1)+\Delta u_{i,0}^{\star}$ are applied for $i=1,\ldots,N$, and the entire procedure is repeated at $t+1$.
If a variational equilibrium is requested, the additional equalities enforcing common Lagrange multipliers for the shared output constraints in (18) are imposed. In Section 6.5, we will show an example of noncooperative MPC and show that it can deviate significantly from the cooperative and centralized MPC solution in which all moves are optimized jointly by solving a quadratic programming (QP) problem with decision variables $\Delta u=\mathop{\rm col}\nolimits(\Delta u_{1},\ldots,\Delta u_{N})$, $\epsilon=\mathop{\rm col}\nolimits(\epsilon_{1},\ldots,\epsilon_{N})$ and cost function given by the sum of all agents’ costs in (18).
@@ 6.1 Linear-quadratic game
$\begin{bmatrix}-0.4&-0.1&-2.1&1.6&-1.8&-0.8\\ 0.5&-1.2&-1.1&-0.9&0.6&2.3\\ 0&-1.1&0.5&-0.6&0.0&1.2\\ -0.7&0&-0.9&-0.2&0.3&-1\end{bmatrix}x\leq\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}$
@@
in $233.9$ ms (11 LM iterations) starting from the initial guess $x^{(0)}=0$. Note that infinitely many different GNEs can exist for the same combination of active constraints, as clearly shown in [23].
Next, we further compare the MILP vs LM approaches on random GNE problems of increasing size. We consider linear-quadratic games with $N$ agents, each controlling $n_{i}=2$ decision variables, with cost functions $f$ as in (19) and shared constraints defined by $n_{g}=2N$ random linear inequalities with unit right-hand side. The results are shown in Figure 1, where we report the CPU time for computing a GNE via MILP vs LM for increasing number $N$ of agents. Computing GNEs on LQ games via MILP is about two orders of magnitude more efficient than by LM with HiGHS, and even more by using Gurobi’s the state-of-the-art MILP solver.
Figure 1: CPU time for computing a GNE via MILP vs LM for increasing number $N$ of agents.
@@ 6.3 Stackelberg game
Here $p=\mathop{\rm col}\nolimits(p_{1,1},\ldots,p_{1,N},p_{2})$, with $p_{1,i}\in[p_{1,\min},p_{1,\max}]$, $p_{2}\in[p_{2,\min},p_{2,\max}]$, is the vector of game parameters chosen by the leader to minimize the loss function
$J(x,p)=\left(\sum_{i=1}^{N}x-D\right)^{2}+\eta\sum_{i=1}^{N}(p_{1,i}-\bar{p}_{1,i})^{2}-\rho\sum_{i=1}^{N}\pi_{i}(x,p)x_{i}.$
The numerical values are set as follows: $N=8$, $C=1$, $D=0.9$, $p_{1,\min}=-5$, $p_{1,\max}=0$, $p_{2,\min}=0$, $p_{2,\max}=0.2$, $\bar{p}_{1,i}=-2$, $\eta=0.1$, $\rho=0.3$, and
$\Gamma=\begin{bmatrix}0&0.2&0&0&0.1&0&0&0\\ 0.2&0&0.1&0&0&0&0&0\\ 0&0.1&0&0.15&0&0&0&0\\ 0&0&0.15&0&0.1&0&0&0\\ 0.1&0&0&0.1&0&0.2&0&0\\ 0&0&0&0&0.2&0&0.1&0\\ 0&0&0&0&0&0.1&0&0.2\\ 0&0&0&0&0&0&0.2&0\end{bmatrix},\qquad a=\begin{bmatrix}1.0\\ 1.5\\ 0.8\\ 1.2\\ 2.0\\ 0.9\\ 1.8\\ 1.1\end{bmatrix}$
Problem (8) is solved by setting $\rho=10^{8}$ with 83 LM iterations in 1.057 s starting from the initial guess $p_{1}^{(0)}=\frac{p_{1,\min}+p_{1,\max}}{2}$, $p_{2}^{(0)}=\frac{p_{2,\min}+p_{2,\max}}{2}$, and $x_{i}^{(0)}=\frac{C}{N}$. The optimal parameters and related GNE are
@@ 6.4 LQR game
The problem is solved as in (6) (without parameter $p_{0}$) in 6.92 s (7 LM iterations), leading to an asymptotically closed-loop matrix $A-BK^{\star}$ with spectral radius 0.7525. For comparison, the centralized (cooperative) LQR solution leads to a spectral radius of 0.3965.
@@ 6.5 Game-theoretic MPC
We consider a linear system as in (16) with $n_{x}=6$ states and $n_{u}=3$ inputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,2,3$, $N=3$ agents), with $A,B\in\mathbb{R}^{6\times 3}$ randomly generated and scaled so that $A$ is stable with spectral radius equal to 0.95. The output matrix is $C\in\mathbb{R}^{3\times 6}$ scaled so that the steady-state gain from inputs to outputs is the identity matrix. Each agent $i$ has state-weighting matrix $Q_{i}=C^{\top}Q_{yi}C$, with $Q_{yi}=e_{i}e_{i}^{\top}$, $Q_{yi}\in\mathbb{R}^{3\times 3}$, and input-increment weighting $Q_{\Delta u,i}=0.5$ for $i=1,2,3$. The prediction horizon is $T=10$ and the constraint horizon is $T_{c}=3$. Input constraints are $0\leq u_{i}\leq 4$ for all $i=1,2,3$, and output constraints are $0\leq y_{i}\leq 5$ for all $i=1,2,3$. The slack penalty is $q_{\epsilon,i}=10^{3}$ for all $i=1,2,3$.
@@
The CPU time for solving each GNEP ranges between $4.76$ and $87.76$ ms (MILP), while the CPU time for solving each centralized MPC problem ranges between $0.42$ and $1.50$ ms (QP).
@@ 6.6 Sparse GNE computations
Hence, if for every agent pair $k$, the corresponding variables $x_{2k-1}=x_{2k}$, we have a Nash equilibrium $x^{\star}\in\mathbb{R}^{N}$. Among the infinitely many Nash equilibria, we want to minimize the game-design objective
@@
where the reference profile $x^{\mathrm{ref}}_{i}=\frac{\lceil(i+1)/2\rceil}{10}$, $i=1,\ldots,N$, and $\alpha_{1}>0$ promotes sparsity in the equilibrium solution. Note that in this case there are no game parameters and $p$ is omitted.
@@
Figure 5 shows the CPU time required to compute the sparse GNE solution as a function of the number $N$ of agents for $\alpha_{1}=2$. The CPU time grows roughly linearly with $N$.

### v2

@@ 4.1 Extracting multiple equilibria
4.2 Variational GNE
As for the nonlinear case, a vGNE can be obtained by enforcing that the Lagrange multipliers associated with the shared constraints are the same for all players, which further reduces the number of variables in (14).
4.2.1 Proximal ADMM solver
For variational LQ games with fixed parameter $p=p_{0}$ and no game objective, we also consider the proximal-ADMM algorithm in [7] as an alternative approach to compute a vGNE. To apply the method, that only handles equality constraints, we reformulate the problem as the following LQ game with $N+1$ players:
$x_{i}^{\star}\in\arg\min_{x_{i}}$ $\frac{1}{2}x^{\prime}Q^{i}x+(c^{i})^{T}x$ (16)
s.t. $\bar{A}x+s=\bar{b},\quad A_{\mathrm{eq}}x=b_{\mathrm{eq}}$
$x_{-i}=x_{-i}^{\star},s=s^{\star},\quad i=1,\ldots,N$
$s^{\star}\in\arg\min_{s}$ $\frac{1}{2}\left\|\bar{A}x^{\star}+s-\bar{b}\right\|_{2}^{2}$
s.t. $\bar{A}x^{\star}+s-\bar{b},s\geq 0$
where, without loss of generality, we have taken $p_{0}=0$. The proximal-ADMM iterations for the reformulation (16) are as follows:
$x_{i}^{k+1}$ $\in\arg\min_{x_{i}}\frac{1}{2}x_{i}^{\prime}Q^{i}_{ii}x_{i}+(c^{i}_{i}+Q^{i}_{i,-i}\bar{x}_{-i}^{k}+(\lambda^{k})^{\top}\bar{A}_{i}+(\mu^{k})^{\top}A_{\mathrm{eq},i})x_{i}+\frac{\gamma}{2}\|x_{i}-x_{i}^{k}\|_{2}^{2}$
$+\frac{\rho}{2}(\|\bar{A}_{i}x_{i}^{k}+\bar{A}_{-i}\bar{x}_{-i}^{k}+s^{k}-\bar{b}\|_{2}^{2}+\|A_{\mathrm{eq},i}x_{i}+A_{\mathrm{eq},-i}\bar{x}_{-i}^{k}-d\|_{2}^{2})$ (17a)
$s^{k+1}$ $\in\arg\min_{s\geq 0}\frac{1+\rho}{2}\left\|\bar{A}x^{k+1}+s-\bar{b}\right\|_{2}^{2}+(\lambda^{k})^{\top}s+\frac{\gamma}{2}\|s-s^{k}\|_{2}^{2}$ (17b)
$\lambda^{k+1}$ $=\lambda^{k}+\rho(\bar{A}x^{k+1}+s^{k+1}-\bar{b})$ (17c)
$\mu^{k+1}$ $=\mu^{k}+\rho(A_{\mathrm{eq}}x^{k+1}-b_{\mathrm{eq}})$ (17d)
where $\bar{x}^{k}_{-i}$ collects the components of $x_{1}^{k+1},\ldots,x_{i-1}^{k+1},x_{i+1}^{k},\ldots,x_{N}^{k}$, $\gamma>0$ is the proximal parameter, and $\rho$ is the augmented Lagrangian parameter. The iterations (17) admit a closed-form solution: the $x_{i}$-update in (17a) is an unconstrained strongly convex quadratic program (QP) that can be solved by caching the Cholesky decompositions of $Q^{i}_{ii}+\gamma I+\rho(\bar{A}_{i}^{\top}\bar{A}_{i}+A_{\mathrm{eq},i}^{\top}A_{\mathrm{eq},i})$ for each player $i$ and then, at each iteration $k$, solving two triangular linear systems; the evaluation of $s^{k+1}$ also admits the closed-form solution
$s^{k+1}=\frac{1}{1+\rho+\gamma}\max\left((1+\rho)(\bar{b}-\bar{A}x^{k+1})+\gamma s^{k}-\mu^{k},0\right)$
as, by completing the squares, (17b) is equivalent to the projection of a vector onto the nonnegative orthant.
@@ 5.1 Game-theoretic linear quadratic regulation
We consider a non-cooperative multi-agent linear-quadratic regulator (LQR) problem where $N$ agents control a shared discrete-time linear dynamical system
@@
where $K_{i}\in\mathbb{R}^{n_{i}\times n_{x}}$ is the LQR gain corresponding to weight matrices $Q_{i}\in\mathbb{R}^{n_{x}\times n_{x}}\succeq 0$, such as $Q_{i}=C^{\top}Q_{yi}C$ with $Q_{yi}\succeq 0$, and $R_{i}\in\mathbb{R}^{n_{i}}$, $R_{i}\succ 0$. We approximate the desired LQR gain $K_{i}$ as a function of the other agents’ gains $K_{-i}$ by running the following Riccati-based iterations
$K_{i}^{k}$ $=(R_{i}+B_{i}^{T}X_{i}^{k}B_{i})^{-1}B_{i}^{T}X_{i}^{k}(A_{i}-B_{-i}K_{-i})$ (19)
$A_{k}$ $=A_{i}-B_{-i}K_{-i}-B_{i}K_{i}^{k}$
$X_{i}^{k+1}$ $=Q_{i}+A_{k}^{T}X_{i}^{k}A_{k}+(K_{i}^{k})^{T}R_{i}K_{i}^{k}$
from $X_{i}^{0}=Q_{i}$ for $k=1,\ldots,N_{\rm LQR}$. Then, we define each agent’s cost function as the squared Frobenius norm of the deviation from its optimal gain $K_{i}^{N_{\rm LQR}}(K_{-i})$:
$f_{i}(K_{i},K_{-i})=\|K_{i}^{N_{\rm LQR}}(K_{-i})-K_{i}\|_{F}^{2}.$
@@
The equilibrium can be computed as the solution of a Nash equilibrium problem by using the nonlinear least-squares approach (6) (without parameter $p_{0}$), using the centralized LQR solution $K_{\rm LQR}$ associated with $(A,B,\sum_{i}Q_{i},\mathrm{blkdiag}(R_{1},\dots,R_{N}))$ as initial guess.
Alternatively, the coupled discrete-time algebraic Riccati equations can be solved using Riccati-based iterations as described in [33, Section III.B] until convergence is reached within a given tolerance.
@@ 5.2 Game-theoretic model predictive control
In (20), $Q_{y,i}\succeq 0$, $Q_{\Delta u,i}\succ 0$, and $\epsilon_{i}\geq 0$ is a slack variable used to soften shared output constraints (with linear penalty $q_{\epsilon,i}\geq 0$). For example, if an agent is interested in controlling only a subset $I_{i}$ of the components of the output vector with unit weights, we have $Q_{y,i}=\sum_{j\in I_{i}}e_{j}e_{j}^{\top}$.
The predicted trajectories $\{x_{k},y_{k},u_{k}\}$ in (20) satisfy the dynamic equations (18) over the prediction horizon $k=0,\ldots,T$, with initial condition $x_{0}=x(t)$ and $u_{-1}=u(t-1)$. Note that each agent optimizes the sequence of input increments $\Delta u_{i}$ and the slack variable $\epsilon_{i}$, so that constraints on inputs, input increments, and slacks are local to each agent, while output constraints are shared among all agents, each one softening them with possibly different slacks $\epsilon_{i}$. Hence, the resulting game-theoretic MPC problem is the GNE problem of finding sequences $\{\Delta u_{i}^{\star}\}_{i=1}^{N}$ (and slacks $\{\epsilon_{i}^{\star}\}_{i=1}^{N}$) optimizing (20) for all agents $i=1,\ldots,N$ simultaneously. In a receding-horizon fashion, only the first inputs $u_{i}(t)=u_{i}(t-1)+\Delta u_{i,0}^{\star}$ are applied for $i=1,\ldots,N$, and the entire procedure is repeated at $t+1$.
The number of constraints in each agent’s problem (20) can be reduced by imposing the constraints only on a shorter constraint horizon of $T_{c}<T$ steps. This can be especially useful when $T$ is large but we want to limit the number of Lagrange multipliers and binary variables involved in the MILP reformulation (14) of the GNEP, in which we remove $p$, $\sigma$, and the objective function (12a), as we are only interested in computing the GNE of the noncooperative MPC problem.
If a variational equilibrium is requested, the additional equalities enforcing common Lagrange multipliers for the shared output constraints in (20) can be imposed. In Section 6.5, we will consider an example of noncooperative MPC and show that it can deviate significantly from the cooperative and centralized MPC solution in which all moves are optimized jointly by solving a quadratic programming (QP) problem with decision variables $\Delta u=\mathop{\rm col}\nolimits(\Delta u_{1},\ldots,\Delta u_{N})$, $\epsilon=\mathop{\rm col}\nolimits(\epsilon_{1},\ldots,\epsilon_{N})$ and cost function given by the sum of all agents’ costs in (20).
@@ 6.1 Linear-quadratic game
$\begin{bmatrix}-0.4&-0.1&-2.1&1.6&-1.8&-0.8\\ 0.5&-1.2&-1.1&-0.9&0.6&2.3\\ 0&-1.1&0.5&-0.6&0.0&1.2\\ -0.7&0&-0.9&-0.2&0.3&-1\end{bmatrix}x\leq\begin{bmatrix}1\\ 1\\ 1\\ 1\end{bmatrix}$.
@@
in $233.9$ ms (11 LM iterations) starting from the initial guess $x^{0}=0$. Note that infinitely many different GNEs can exist for the same combination of active constraints, as clearly shown in [24]. The same vGNE is found by running the proximal ADMM approach described in Section 4.2.1, where we set $\gamma=1$, $\rho=1$, $x^{0}=0$, $\lambda^{0}=0$, which takes $14.6$ ms (229 ADMM iterations).
Next, we further compare the MILP, LM, and proximal ADMM approaches on random GNE problems of increasing size. We consider linear-quadratic games with $N$ agents, each controlling $n_{i}=2$ decision variables, with cost functions $f$ as in (21) and shared constraints defined by $n_{g}=2N$ random linear inequalities with unit right-hand side. The results are shown in Figure 1(a), where, for an increasing number $N$ of agents, we report the CPU time for computing a GNE via MILP and for computing a vGNE by LM and proximal ADMM. Computing GNEs on LQ games via MILP with HiGHS is about two orders of magnitude more efficient than by LM, and even more by using Gurobi’s the state-of-the-art MILP solver; it also outperforms proximal ADMM (which takes between 258 and 1794 iterations). Figure 1(b) shows instead the CPU time taken when fixing $N=5$ agents and increasing the number of shared constraints $n_{g}$ up to 300, showing that MILP methods degrade more in performance compared to proximal ADMM due to the increasing number of binary variables in (14).
(a) CPU time vs number $N$ of agents, under $n_{g}=2N$ shared constraints.
(b) CPU time vs number of shared constraints, with $N=5$ agents.
Figure 1: CPU time for computing a GNE via MILP vs LM for increasing number of agents and constraints.
@@ 6.3 Stackelberg game
Here $p=\mathop{\rm col}\nolimits(p_{1,1},\ldots,p_{1,N},p_{2})$ is the vector of game parameters chosen by the leader to minimize the loss function
$J(x,p)=\left(\sum_{i=1}^{N}x-D\right)^{2}+\eta\sum_{i=1}^{N}(p_{1,i}-\bar{p}_{1,i})^{2}-\rho\sum_{i=1}^{N}\pi_{i}(x,p)x_{i}$
with $p_{1,i}\in[p_{1,\min},p_{1,\max}]$ and $p_{2}\in[p_{2,\min},p_{2,\max}]$. The numerical values are set as follows: $N=8$, $C=1$, $D=0.9$, $p_{1,\min}=-5$, $p_{1,\max}=0$, $p_{2,\min}=0$, $p_{2,\max}=0.2$, $\bar{p}_{1,i}=-2$, $\eta=0.1$, $\rho=0.3$, and
$\Gamma=\begin{bmatrix}0&0.2&0&0&0.1&0&0&0\\ 0.2&0&0.1&0&0&0&0&0\\ 0&0.1&0&0.15&0&0&0&0\\ 0&0&0.15&0&0.1&0&0&0\\ 0.1&0&0&0.1&0&0.2&0&0\\ 0&0&0&0&0.2&0&0.1&0\\ 0&0&0&0&0&0.1&0&0.2\\ 0&0&0&0&0&0&0.2&0\end{bmatrix}\!\!,\qquad a=\begin{bmatrix}1.0\\ 1.5\\ 0.8\\ 1.2\\ 2.0\\ 0.9\\ 1.8\\ 1.1\end{bmatrix}\!\!$.
By setting $\rho=10^{8}$, (8) is solved with 83 LM iterations in 1.057 s starting from the initial guess $p_{1}^{(0)}=\frac{p_{1,\min}+p_{1,\max}}{2}$, $p_{2}^{(0)}=\frac{p_{2,\min}+p_{2,\max}}{2}$, and $x_{i}^{(0)}=\frac{C}{N}$. The optimal parameters and related GNE are
@@ 6.4 LQR game
The problem is solved as in (6) (without parameter $p_{0}$) in 6.05 s (6 LM iterations), leading to an asymptotically closed-loop matrix $A-BK^{\star}$ with spectral radius 0.7525. For comparison, the centralized (cooperative) LQR solution leads to a spectral radius of 0.3965. Solving the same problem by using the Riccati-based iterations described in [33, Section III.B] leads to the same solution in 0.11 s (7 Riccati iterations).
@@ 6.5 Game-theoretic MPC
We consider a linear system as in (18) with $n_{x}=6$ states, $n_{u}=3$ inputs, and $n_{y}=3$ outputs, each one controlled by a different agent ($n_{i}=1$ for all $i=1,2,3$, $N=3$ agents), with $A,B\in\mathbb{R}^{6\times 3}$ randomly generated and scaled so that $A$ is stable with spectral radius equal to 0.95. The output matrix $C\in\mathbb{R}^{3\times 6}$ is scaled so that the steady-state gain from inputs to outputs is the identity matrix. Each agent $i$ has state-weighting matrix $Q_{i}=C^{\top}Q_{y,i}C$, with $Q_{y,i}=e_{i}e_{i}^{\top}$, $Q_{y,i}\in\mathbb{R}^{3\times 3}$, and input-increment weighting $Q_{\Delta u,i}=0.5$ for $i=1,2,3$. The prediction horizon is $T=10$ and the constraint horizon is $T_{c}=3$. Input constraints are $0\leq u_{i}\leq 4$ for all $i=1,2,3$, and output constraints are $0\leq y_{i}\leq 5$ for all $i=1,2,3$. The slack penalty is $q_{\epsilon,i}=10^{3}$ for all $i=1,2,3$.
@@
The CPU time for solving each GNEP ranges between $4.76$ and $87.76$ ms (HiGHS MILP), while the CPU time for solving each centralized MPC problem ranges between $0.42$ and $1.50$ ms (QP).
@@ 6.6 Sparse GNE computations
Clearly, if for every agent-pair $k$ the corresponding variables $x_{2k-1}=x_{2k}$, we have a Nash equilibrium $x^{\star}\in\mathbb{R}^{N}$. Among the infinitely many Nash equilibria, we want to minimize the game-design objective
@@
where the components of the reference vector $x^{\mathrm{ref}}_{i}=\frac{\lceil(i+1)/2\rceil}{10}$, $i=1,\ldots,N$, and $\alpha_{1}>0$ promotes sparsity in the equilibrium solution. Note that in this case there are no game parameters and $p$ is omitted.
@@
Figure 5 shows the CPU time required to compute the sparse GNE solution as a function of the number $N$ of agents for $\alpha_{1}=2$ using L-BFGS-B. The CPU time grows roughly linearly with $N$.